The letters of the word Mississippi are placed in a hat. What is the probability that the first two letters drawn are both i if the first letter drawn is not replaced?
Answer by nerdybill(7009)

The letters of the word Mississippi are placed in a hat. What is the probability that the first two letters drawn are both i if the first letter drawn is not replaced?
.
Total letters is 11
Number of i's is 4
.
probability of drawing the first i is 4/11
probability of drawing the second i is 3/10
.
4/11 * 3/10 = 12/110 = 0.11